OpenStudy (anonymous):

solve the following system of equations with steps shown: F - 9/16FL - 3/4F1 - F2 F - 3/4FL - F1 + F2 F - FL + F1 + 3/4F2

OpenStudy (anonymous):

plz dear god help me

OpenStudy (anonymous):

i do not remember how to do this stuff

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i know F = 241..that's given

OpenStudy (anonymous):

i have to solve for FL F1 AND F2

OpenStudy (anonymous):

- 9/16FL - 3/4F1 - F2 =-241 - 3/4FL - F1 + F2 =-241 - FL + F1 + 3/4F2=-241

OpenStudy (anonymous):

just do matrix A x= B x= A^-1 B

OpenStudy (anonymous):

i don't know how to do matrix

OpenStudy (anonymous):

i learned really basic stuff with system of equations..so this stuff is above my head

OpenStudy (anonymous):

oh, I see,

OpenStudy (anonymous):

you got 3 equation, let's reduce it - 9/16FL - 3/4F1 - F2 =-241 - 3/4FL - F1 + F2 =-241 - FL + F1 + 3/4F2=-241 F1=( -241 +FL- 3/4 F2) plug in that everytime you see F1 - 9/16FL - 3/4F1 - F2 =-241 - 3/4FL - F1 + F2 =-241 - 9/16FL - 3/4( -241 +FL- 3/4 F2) - F2 =-241 - 3/4FL - ( -241 +FL- 3/4 F2) + F2 =-241

OpenStudy (anonymous):

now you got two by two equation

OpenStudy (anonymous):

umm...if i give you a dollar will u do the rest? lol

OpenStudy (anonymous):

I can give you answer, fine?

OpenStudy (anonymous):

yea..that'd be a great help

OpenStudy (anonymous):

i just need to write down the steps to it

OpenStudy (anonymous):

hold on , you variable names don't help

OpenStudy (anonymous):

not sure if this helps...but the answers should come close to FL = 5F/3, F1=F/2, F2 = F/3

OpenStudy (anonymous):

was the answer to a similar problem we had with different numbers...i couldn't solve that one either

OpenStudy (anonymous):

\[FL\to \frac{9 F}{7},F1\to \frac{5 F}{28},F2\to \frac{F}{7}\]

OpenStudy (anonymous):

so so cool...can u show me the steps to get that?

OpenStudy (anonymous):

yep, but I will tell you method called Inverse Soloution 1) Write you system as matrix I send number to other sides \[\left( \begin{array}{ccc} -\frac{9}{16} & -\frac{3}{4} & -1 \\ -\frac{3}{4} & -1 & 1 \\ -1 & 1 & \frac{3}{4} \end{array} \right).\left( \begin{array}{c} \text{FL} \\ \text{F1} \\ \text{F2} \end{array} \right)=\left( \begin{array}{c} -241 \\ -241 \\ -241 \end{array} \right)\] A x = B we want to find x, so we multiply inverse of A to both side , leaving only x on left side x= A^-1 B A^-1=\[\left( \begin{array}{ccc} -\frac{4}{7} & -\frac{1}{7} & -\frac{4}{7} \\ -\frac{1}{7} & -\frac{13}{28} & \frac{3}{7} \\ -\frac{4}{7} & \frac{3}{7} & 0 \end{array} \right)\] \[\left( \begin{array}{c} \text{FL} \\ \text{F1} \\ \text{F2} \end{array} \right)=\left( \begin{array}{ccc} -\frac{4}{7} & -\frac{1}{7} & -\frac{4}{7} \\ -\frac{1}{7} & -\frac{13}{28} & \frac{3}{7} \\ -\frac{4}{7} & \frac{3}{7} & 0 \end{array} \right).\left( \begin{array}{c} -241 \\ -241 \\ -241 \end{array} \right)\]

OpenStudy (anonymous):

and that's it?

OpenStudy (anonymous):

when we multiply these two matrix we should get our answer

OpenStudy (anonymous):

thank you so very much...i owe you huge

OpenStudy (anonymous):

I assume this is for your science class?

OpenStudy (anonymous):

physics of meteorology yes

OpenStudy (anonymous):

yeah,you would use matrix for solving equations in application

OpenStudy (anonymous):

i starred that

OpenStudy (anonymous):

you started that?

OpenStudy (anonymous):

no...i starred it..as in saved it for future reference

OpenStudy (anonymous):

oh yeah, if you have graphing calculator (Like TI-83,84) you can also do it there